# Learning from Natural Language Explanations for Generalizable Entity Matching

### v1

@@ 4.1 Deep learning in Entity Resolution
Other significant contributions include Seq2SeqMatcher Nie et al. (2019), focusing on sequence-to-sequence matching, and HierMatcher Fu et al. (2021), which adopts a hierarchical approach. The use of pre-trained language models has also gained traction, as evidenced by methods such as R-SupCon, Ditto, Rotom, and Sudowoodo, discussed in various studies Brunner and Stockinger (2020); Peeters et al. (2020); Li et al. (2021); Miao et al. (2021); Wang et al. (2023b). These methods collectively represent the cutting-edge techniques in the realm of entity matching.
@@ 4.2 Reasoning in LLMs
Most recently, Entity Matching via LLMs has shown promising results Peeters and Bizer (2023c, b). In these works, both zero-shot and fine-tuning approaches have been explored. Beyond entity matching, in-context learning (ICL) with LLMs has become a dominant strategy, enabling these models to perform tasks with task conditioning and minimal task demonstrations Brown et al. (2020); Xie et al. (2021). This approach has demonstrated strong performance Zhao et al. (2021); Liu et al. (2021) and streamlined experimentation with LLMs, as it eliminates the need for model training. However, the adoption of ICL has highlighted the sensitivity of LLMs to prompt selection Lu et al. (2021); Margatina et al. (2023), making prompt engineering for various tasks a challenging and time-consuming process. Nonetheless, data-driven signals, such as selecting semantically similar demonstrations using text retrievers, have proven to be effective Lu et al. (2021); Margatina et al. (2023), offering a more systematic approach to prompt engineering.
@@ Appendix D OOD Performance in Neural Entity Matching
Alpaca (7B) Mistral-7B-Instruct Falcon-Instruct (7B) FlanT5-XXL (11B) Flan-UL2 (20B)
@@
iTunes-Amazon 15.61 95.66 26.84 28.32 87.59 42.80 11.57 98.47 20.71 15.46 77.77 25.79 20.69 85.12 33.29
WDC-Computers 29.74 84.24 43.96 32.49 64.76 43.27 29.59 91.20 44.68 23.71 82.45 36.83 92.55 60.41 73.10
WDC-Cameras 30.57 85.40 45.02 33.08 72.24 45.31 26.99 90.16 41.54 36.05 87.77 51.11 80.51 61.97 70.03
WDC-Watches 35.49 85.36 50.14 34.47 75.68 47.37 11.17 83.18 19.70 34.19 85.44 48.84 84.13 68.82 75.71
WDC-Shoes 32.79 62.24 42.95 32.51 78.35 51.64 36.43 75.19 49.08 29.22 65.09 29.22 75.48 50.17 60.28

### v2

@@ 4.1 Deep learning in Entity Resolution
Other significant contributions include Seq2SeqMatcher Nie et al. (2019); Wang and Zhang (2024), focusing on sequence-to-sequence matching, and HierMatcher Fu et al. (2021), which adopts a hierarchical approach. The use of pre-trained language models has also gained traction, as evidenced by methods such as R-SupCon, Ditto, Rotom, and Sudowoodo, discussed in various studies Brunner and Stockinger (2020); Peeters et al. (2020); Li et al. (2021); Miao et al. (2021); Wang et al. (2023b, 2024); Zeakis et al. (2023); Genossar et al. (2023). These methods collectively represent the cutting-edge techniques in the realm of entity matching.
@@ 4.2 Reasoning in LLMs
Most recently, Entity Matching via LLMs has shown promising results Peeters and Bizer (2023c, b); Fan et al. (2024). In these works, both zero-shot and fine-tuning approaches have been explored. Beyond entity matching, in-context learning (ICL) with LLMs has become a dominant strategy, enabling these models to perform tasks with task conditioning and minimal task demonstrations Brown et al. (2020); Xie et al. (2021). This approach has demonstrated strong performance Zhao et al. (2021); Liu et al. (2021) and streamlined experimentation with LLMs, as it eliminates the need for model training. However, the adoption of ICL has highlighted the sensitivity of LLMs to prompt selection Lu et al. (2021); Margatina et al. (2023), making prompt engineering for various tasks a challenging and time-consuming process. Nonetheless, data-driven signals, such as selecting semantically similar demonstrations using text retrievers, have proven to be effective Lu et al. (2021); Margatina et al. (2023), offering a more systematic approach to prompt engineering.
@@ Appendix D OOD Performance in Neural Entity Matching
Alpaca (7B) Mistral-7B-Ins Falcon-Ins (7B) FlanT5-XXL Flan-UL2
@@
iT-A 15.61 95.66 26.84 28.32 87.59 42.80 11.57 98.47 20.71 15.46 77.77 25.79 20.69 85.12 33.29
W-Com 29.74 84.24 43.96 32.49 64.76 43.27 29.59 91.20 44.68 23.71 82.45 36.83 92.55 60.41 73.10
W-Cam 30.57 85.40 45.02 33.08 72.24 45.31 26.99 90.16 41.54 36.05 87.77 51.11 80.51 61.97 70.03
W-Wat 35.49 85.36 50.14 34.47 75.68 47.37 11.17 83.18 19.70 34.19 85.44 48.84 84.13 68.82 75.71
W-Sh 32.79 62.24 42.95 32.51 78.35 51.64 36.43 75.19 49.08 29.22 65.09 29.22 75.48 50.17 60.28